# Echoes of Socratic Doubt: Embracing Uncertainty in Calibrated Evidential Reinforcement Learning

### v1

@@ Abstract
We present a novel statistical approach to incorporating uncertainty awareness in model-free distributional reinforcement learning involving quantile regression-based deep Q networks. The proposed algorithm, Calibrated Evidential Quantile Regression in Deep Q Networks (CEQR-DQN), aims to address key challenges associated with separately estimating aleatoric and epistemic uncertainty in stochastic environments. It combines deep evidential learning with quantile calibration based on principles of conformal inference to provide explicit, sample-free computations of global uncertainty as opposed to local estimates based on simple variance, overcoming limitations of traditional methods in computational and statistical efficiency and handling of out-of-distribution (OOD) observations. Tested on a suite of miniaturized Atari games (i.e., MinAtar), CEQR-DQN is shown to surpass similar existing frameworks in scores and learning speed. Its ability to rigorously evaluate uncertainty improves exploration strategies and can serve as a blueprint for other algorithms requiring uncertainty awareness.
@@ 1 Introduction
Deep reinforcement learning (DRL) is a subset of machine learning (ML) under the umbrella of AI that involves using neural networks to solve sequential decision-making problems that require interaction between an agent and an environment. Recent applications of DRL include drone racing [Kaufmann et al., 2023], healthcare diagnosis and treatment [Yu et al., 2023], and chip design [Mirhoseini et al., 2021], amongst many others. There are several types of DRL algorithms, although in this study we focus on deep Q networks with discrete action spaces [Mnih et al., 2015]. In regards to processing data, a key difference between most DRL algorithms and other ML algorithms is their foundation in the Markov Decision Process (MDP) [Puterman, 1994], a mathematical framework for discrete-time decision-making when the desired outcome (e.g., maximum cumulative reward) is stochastic yet partially amenable to a user. Accordingly, DRL is susceptible to uncertainty, which is reflected directly in each interaction and the end result. However, due to the nature of the algorithm’s interactive data processing, one is given explicit parametric control over whether to exploit or explore the environment when seeking rewards, which can be manipulated with uncertainty awareness. Unlike most supervised ML tasks, where uncertainty awareness has seen the most interest lately, reinforcement learning can directly and explicitly incorporate reliable uncertainty measures as feedback in the exploitation/exploration trade-off and choosing actions.
@@
  • Combining calibrated quantile regression with deep evidential learning enhances the algorithm’s uncertainty awareness by providing global network uncertainty estimates as opposed to local estimates based solely on the variance and expectation of Q-value quantiles, which can be biased. As a result, the proposed CEQR-DQN algorithm achieves meets or exceeds scores on benchmarks such as MinAtar [Young and Tian, 2019] set by similar previous work with notably faster learning.
@@ 3 Uncertainty-Aware Reinforcement Learning
     $u^{5^{th}}_{evi}=[|(\gamma^{5^{th}}_{evi}+\sigma_{evi})-(\gamma^{5^{th}}_{evi}-\sigma_{evi})|]$
@@
     $u^{95^{th}}_{evi}=[|(\gamma^{95^{th}}_{evi}+\sigma_{evi})-(\gamma^{95^{th}}_{evi}-\sigma_{evi})|]$
@@ 3.1 Deep Evidential Learning for Action Choice
$\begin{split}\mathcal{L}_{reg}&=\mathbb{I}\{y\geq\hat{y}\}q(y-\hat{y})\\ &+\mathbb{I}\{y<\hat{y}\}(1-q)(\hat{y}-y)\text{,}\\ \end{split}$ (6)
@@
A visual example of these quantities estimated within a relatively complex synthetic dataset (consisting of $1000$ unseen samples) using calibrated quantile regression is provided in Fig. 2. It is shown that while the predicted quantiles are well calibrated around the noisy data, the underlying uncertainty associated with estimating them is appropriately moderate in the range $[-3,3]$ where the model was trained, and highly inflated for OOD samples outside this range. This example represents the complementary relationship of calibrated quantile regression and evidential learning in accurately covering a complicated distribution of targets while clearly indicating a lack of confidence, especially in samples the model had not seen before. In Section 3.2, we discuss the methods for the calibration and how it is applied to the predicted target distribution and evidential quantiles in DRL.
@@ 3.2 Quantile Calibration via Conformal Methods
[…]
For the evidential sample mean $\gamma$ we only predict and calibrate the $5^{th}$ and $95^{th}$ percentiles so as to establish a $90\%$ evidence-based confidence interval $C_{evi}(X)$ around the TD target from which we gauge uncertainty. However, because the computed TD target is of size $N$ quantiles, we must estimate $N$ quantiles of $\gamma$ for each percentile level. To establish the confidence interval as is done in [Stutts et al., 2023], we reuse $\mathcal{L}_{cal}$ except that $q_{i}$ in the indicator functions is replaced with the desired marginal coverage rate $p=0.9$, ensuring that the focus of the calibration is on marginal coverage of the TD target over each training batch sampled from experience replay. Additionally, we seek to empirically center the TD target within the computed $C_{evi}(X)$ interval so as to enhance its accuracy. To do this, we add another loss function
$\begin{split}\mathcal{L}_{interval}=\frac{1}{N}\sum_{i=1}^{N}\leavevmode\nobreak\ &(\hat{q}^{95^{th}}_{i}-\leavevmode\nobreak\ \hat{q}^{5^{th}}_{i})\leavevmode\nobreak\ +\\ &\mathbb{I}\{y<\hat{q}^{5^{th}}_{i}\}\frac{2}{q}(\hat{q}^{5^{th}}_{i}-\leavevmode\nobreak\ y)\leavevmode\nobreak\ +\\ &\mathbb{I}\{y>\hat{q}^{95^{th}}_{i}\}\frac{2}{q}(y-\hat{q}^{95^{th}}_{i})\text{,}\\ \end{split}$ (12)
which is computed twice and averaged for $q=0.05$ and $q=0.95$.
@@
Figure 3: MinAtar Results: Comparison of running average training scores between CEQR-DQN and UA-DQN on the simplified MinAtar games with $2.5$ million frames. Max episode scores for CEQR-DQN in each game are $103$, $4048$, $64$, $119$, and $3095$, while for UA-DQN they are $127$, $29$, $70$, $44$, and $1847$.

### v3

@@ Abstract
We present a novel statistical approach to incorporating uncertainty awareness in model-free distributional reinforcement learning involving quantile regression-based deep Q networks. The proposed algorithm, Calibrated Evidential Quantile Regression in Deep Q Networks (CEQR-DQN), aims to address key challenges associated with separately estimating aleatoric and epistemic uncertainty in stochastic environments. It combines deep evidential learning with quantile calibration based on principles of conformal inference to provide explicit, sample-free computations of global uncertainty as opposed to local estimates based on simple variance. This approach overcomes limitations of traditional methods in computational and statistical efficiency and handling of out-of-distribution (OOD) observations. Tested on a suite of miniaturized Atari games (i.e., MinAtar), CEQR-DQN is shown to surpass similar existing frameworks in scores and learning speed. Its ability to rigorously evaluate uncertainty improves exploration strategies and can serve as a blueprint for other algorithms requiring uncertainty awareness.
@@ 1 Introduction
Deep reinforcement learning (DRL) is a subset of machine learning (ML) under the umbrella of AI that involves using neural networks to solve sequential decision-making problems that require interaction between an agent and an environment. Recent applications of DRL include drone racing [Kaufmann et al., 2023], healthcare diagnosis and treatment [Yu et al., 2023], and chip design [Mirhoseini et al., 2021], which are all distinct tasks. As such, there are several types of DRL algorithms, although in this study we focus on deep Q networks with discrete action spaces [Mnih et al., 2015]. In regards to processing data, a key difference between most DRL algorithms and other ML algorithms is their foundation in the Markov Decision Process (MDP) [Puterman, 1994], a mathematical framework for discrete-time decision-making when the desired outcome (e.g., maximum cumulative reward) is stochastic yet partially amenable to a user. Accordingly, DRL is susceptible to uncertainty, which is reflected directly in each interaction and the end result. However, due to the nature of the algorithm’s interactive data processing, one is given explicit parametric control over whether to exploit or explore the environment when seeking rewards, which can be manipulated with uncertainty awareness. Unlike most supervised ML tasks, where uncertainty awareness has seen the most interest lately, reinforcement learning can directly and explicitly incorporate reliable uncertainty measures as feedback in the exploitation/exploration trade-off and choosing actions.
@@
  • Combining calibrated quantile regression with deep evidential learning enhances the algorithm’s uncertainty awareness by providing global network uncertainty estimates over the entire known sample space as opposed to local estimates based solely on the variance and expectation of Q-value quantiles, which can be biased. As a result, the proposed CEQR-DQN algorithm achieves meets or exceeds scores on benchmarks such as MinAtar [Young and Tian, 2019] set by similar previous work with notably faster learning.
@@ 3 Uncertainty-Aware Reinforcement Learning
     $u^{5^{th}}_{evi}=[|(\gamma^{5^{th}}_{evi}+\sigma^{5^{th}}_{evi})-(\gamma^{5^{th}}_{evi}-\sigma^{5^{th}}_{evi})|]$
@@
     $u^{95^{th}}_{evi}=[|(\gamma^{95^{th}}_{evi}+\sigma^{95^{th}}_{evi})-(\gamma^{95^{th}}_{evi}-\sigma^{95^{th}}_{evi})|]$
@@ 3.1 Deep Evidential Learning for Action Choice
$\begin{split}\mathcal{L}_{reg}=\Phi&[\mathbb{I}\{y\geq\hat{y}\}q(y-\hat{y})\\ &+\mathbb{I}\{y<\hat{y}\}(1-q)(\hat{y}-y)]\text{,}\\ \end{split}$ (6)
@@
A visual example of these quantities estimated within a relatively complex synthetic dataset (consisting of $1000$ unseen samples) using calibrated quantile regression is provided in Fig. 2. It is shown that while the predicted quantiles are well calibrated around the noisy data, the underlying uncertainty associated with estimating them is appropriately moderate in the range $[-3,3]$ where the model was trained, and highly inflated for OOD samples outside this range. Accordingly, the presence of OOD samples hinders the model’s ability to achieve the desired marginal coverage of $90\%$. This example represents the complementary relationship of calibrated quantile regression and evidential learning in accurately covering a complicated distribution of targets while clearly indicating a lack of confidence, especially in samples the model had not seen before. In Section 3.2, we discuss the methods for the calibration and how it is applied to the predicted target distribution and evidential quantiles in DRL.
@@ 3.2 Quantile Calibration via Conformal Methods
$q$ is the target quantile level, $\hat{q}$ is the predicted quantile, $p^{cov}_{avg}$ represents the average probability that $y$ falls under $\hat{q}$, and $\lambda_{cal}$ is a hyperparameter balancing coverage of $y$ (COVobj) with minimizing the distance between $\hat{q}$ and $1-\hat{q}$ in covering $y$ (SHARPobj). This function individually calibrates each quantile predicted for $Z_{\theta}$ such that they are as close to the target levels as possible given success in training.
For the evidential sample mean $\gamma$ we only predict and calibrate the $5^{th}$ and $95^{th}$ percentiles so as to establish a $90\%$ evidence-based confidence interval $C_{evi}(X)$ around the TD target from which we gauge uncertainty. However, because the computed TD target is of size $N$ quantiles, we must estimate $N$ quantiles of $\gamma$ for each percentile level. To establish the confidence interval as is done in [Stutts et al., 2023], we reuse $\mathcal{L}_{cal}$ except that $q_{i}$ in the indicator functions is replaced with the desired marginal coverage rate $p=0.9$, ensuring that the focus of the calibration is on marginal coverage of the TD target over each training batch sampled from experience replay. Additionally, we seek to empirically center the TD target within the computed $C_{evi}(X)$ interval so as to enhance its accuracy. To do this, we add another loss function:
$\begin{split}\mathcal{L}_{interval}=\frac{1}{N}\sum_{i=1}^{N}\leavevmode\nobreak\ &(\hat{q}^{95^{th}}_{i}-\leavevmode\nobreak\ \hat{q}^{5^{th}}_{i})\leavevmode\nobreak\ +\\ &\mathbb{I}\{y<\hat{q}^{5^{th}}_{i}\}\frac{2}{q}(\hat{q}^{5^{th}}_{i}-\leavevmode\nobreak\ y)\leavevmode\nobreak\ +\\ &\mathbb{I}\{y>\hat{q}^{95^{th}}_{i}\}\frac{2}{q}(y-\hat{q}^{95^{th}}_{i})\text{.}\\ \end{split}$ (12)
@@
Figure 3: MinAtar Results: Comparison of second-order running average training scores between CEQR-DQN and UA-DQN across 25 random seeds on the simplified MinAtar games with $2.5$ million frames. Max episode scores for CEQR-DQN in each game are $158$, $4481$, $65$, $199$, and $4677$, while for UA-DQN they are $95$, $2457$, $69$, $95$, and $2231$. Results are also shown for CEQR-DQN without calibration.